OpenStudy (beth12345):

A 2.6kg laboratory cart is given a push and moves with a speed of 2.0m/s toward a solid barrier, where it is brought to rest by a spring bumper. How much energy will be stored in the spring (elastic potential energy) at the moment when the spring is fully compressed? What is the average force exerted by the spring if it is compressed 0.12m? (Why is it necessary to specify average force in this situation?)

OpenStudy (anonymous):

(1) The kinetic energy of the cart. (2) The work done by the spring on the cart is the change in its kinetic energy (Work-Energy Theorem), which you know from (1). You could estimate the average force by \(W=F_{avg}\Delta x\), where \(W\) is the work done by the spring.

OpenStudy (beth12345):

what's the change in x stand for?

OpenStudy (beth12345):

\[\Delta x\]?

OpenStudy (anonymous):

delta x stands for change in the length of the spring!

OpenStudy (anonymous):

were u able to understand?

OpenStudy (beth12345):

not really

OpenStudy (anonymous):

what is the kinetic energy of cart first?

OpenStudy (anonymous):

take the cart and the spring as the system,the forces between them will be internal,net external force on the system is zero so u can use principle of conservation of energy

OpenStudy (anonymous):

initial mechanical energy of the system=final ME

OpenStudy (anonymous):

now try to frame the equation

OpenStudy (beth12345):

\[Ek=1/2 \times g \times h\] \[=1/2 \times 2.6kg \times 2m/s ^{2}\]\[=5.2 J\]

OpenStudy (anonymous):

yes u got it right,now what's the potential energy stored in the spring initially?

OpenStudy (beth12345):

would it be the same? since Ek=Ep?

OpenStudy (anonymous):

do u know the expression for potential e stored in a spring?

OpenStudy (beth12345):

no

OpenStudy (anonymous):

its (1/2)*k*x^2 where k=spring constant of the spring, x=compession or elongatian from natural length

OpenStudy (anonymous):

initially the spring is not elongated or compressed, so what's theEp stored in it initially?

OpenStudy (beth12345):

0?

OpenStudy (anonymous):

one point to be corrected: k.e=1/2*m*v^2 correct that beth

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

mechanical energy= Ep+Ek so what's the initial MEof the system?

OpenStudy (beth12345):

5.2J

OpenStudy (anonymous):

yes now tke the final situation when the spring is fully compressed and calculate the finalME of the system and then apply principle of COE

OpenStudy (beth12345):

i don't really know what the spring constant is

OpenStudy (anonymous):

ok no problem what will happen to the cart when spring is fully comprsd?

OpenStudy (beth12345):

it will bounce back?

OpenStudy (beth12345):

or collect potential energy?

OpenStudy (anonymous):

yes it will,but that's not what we r luking for we r interstd at the moment wen spring is fully cprsd

OpenStudy (anonymous):

it will collect PE only wen it has lost all of its Ek to the spring

OpenStudy (anonymous):

so wat wud be the Ek of the cart at that particular moment?

OpenStudy (beth12345):

oh would it be 0 since the velocity is 0?

OpenStudy (anonymous):

yes exactly now conside the PE of the spring and useCOE to get PE

OpenStudy (anonymous):

the final ME of the system is only due to thePE of the spring now this is what COE says initialME=fialME use this equation

OpenStudy (beth12345):

what does ME and COE stand for again?

OpenStudy (anonymous):

ME=mechanical energy(Ep+Ek) COE=conservation of ME

OpenStudy (beth12345):

would the answer be 5.2J fir the fianal ME

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and as i mentioned earlier finalME is due to the potential energy stored in the spring

OpenStudy (beth12345):

ok, thank you :)

OpenStudy (anonymous):

there is one more way to solve this q and that is by using work energy theorem

OpenStudy (beth12345):

ok, ill try that later but right now I have to catch my bus, thank you so much for helping me out :)

OpenStudy (anonymous):

ur welcome:)